Circular No. 7996

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002gy IN UGC 2701 Further to IAUC 7972, B. Beutler and W. D. Li report the LOTOSS

discovery of an apparent supernova (mag ∼ 17.6) on an unfiltered CCD KAIT image taken on Oct. 16.4 UT. The new object is located at α = 3^{h}22^{m}36^{s}.45, δ = +9^{o}28′22′′.0 (equinox 2000.0), which is 2′′.8 east and 3′′.0 south of the nucleus of UGC 2701. M. Hamuy, Carnegie Institution of Washington, confirmed SN 2002gy on J and H images taken with the Las Campanas Dupont 2.5-m telescope (+ WIRC) on Oct. 17.4. A KAIT image taken on Oct. 9.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002gz Further to IAUC 7995, W. M. Wood-Vasey, G. Aldering, P. Nugent,

and K. Li report the discovery of an apparent supernova (mag 18.3) on unfiltered NEAT images taken with the Palomar 1.2-m Schmidt telescope on Oct. 15.48, 15.49, and 15.50 UT (confirmed on Oct. 18 at mag 18.2). The new object is located at α = 2^{h}34^{m}10^{s}.36, δ = −0^{o}53′18′′.2 (equinox 2000.0), which is 4′′ east and 1′′ south of the center of the apparent host galaxy. No point source is visible at this location in previous NEAT images taken on Sept. 22 (limiting mag 20.2 at S/N = 5).

SUPERNOVA 2002gv G. Miknaitis, University of Washington, reports that a CCD spectrum

of SN 2002gv (cf. IAUC 7994), taken on Oct. 14 UT with the Baade 6.5-m telescope (+ LDSS-2) at the Las Campanas Observatory by K. Krisciunas, N. Morrell, and M. M. Phillips, is consistent with that of a type-Ia supernova, ≈ 10 days after maximum, at z ∼ 0.24. A spectrum of its host galaxy taken on Oct. 15, confirms a redshift of 0.25.

SUPERNOVA 2002gw IN NGC 922 Unfiltered CCD magnitudes: Oct. 11.633 UT, 18.3 (K. Itagaki, Yama-

gata, Japan; prediscovery image; forwarded by S. Nakano, Sumoto, Japan); 17.92, 17.8 (L. A. G. Monard, Pretoria, S. Africa).

POSSIBLE NOVA IN SMALL MAGELLANIC CLOUD Another broadband V CCD magnitude of this object (cf. IAUC 7995)

from W. Liller, Viña del Mar, Chile: Oct. 19.131 UT, 13.19. Visual magnitude estimate by A. Pearce, Nedlands, W. Australia: Oct. 18.654, 13.2.

2002 October 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green